# Entanglement Dynamics in Monitored Systems and the Role of Quantum Jumps

### 20 Dec 2023 v1

@@ IV Entanglment Dynamics under Quantum Jumps
In this section we present our results for the QJ dynamics of the SSH and Ising chain model, focusing first on the dynamics of the entanglement entropy. Unless specified otherwise, throughout this section we consider an initial product state of the form $|\Psi(0)\rangle=|0,1,0,1,\dots,0,1\rangle$ and open boundary conditions. We fix the hopping along the chains as units of energy and inverse time, $J=1$. We integrate the dynamics with a time-step $dt\propto 1/({4\gamma L})$. Throughout this work we consider a system of size $L=128$.
@@ IV.1 Entanglement Dynamics in the Monitored SSH Model
We start discussing the entanglement dynamics for the monitored SSH. In Fig. 3 (a-c), we show the evolution of the entanglement entropy for a sample of quantum trajectories for increasing values of the monitoring rate $\gamma=0.5,1.5,2.5$, corresponding to the volume-law [panels (a,b)] and area-law [panel (c)] phase of the no-click limit, cf. Fig. 2. For comparison, we also plot the average entanglement entropy (obtained averaging over the different trajectories, according to Eq. (12)) as well as the value obtained in the non-Hermitian no-click limit. We see that, for weak monitoring, when the no-click dynamics displays a linear growth corresponding to the volume law phase, the stochastic dynamics due to quantum jumps suppress the entanglement very effectively and bring the system to a stationary state which is less entangled than in the no-click limit. In this regime, an agreement with pure non-Hermitian evolution is present only for short periods of time. We also notice that for weak monitoring, trajectory-to-trajectory fluctuations are rather broad, and certain outliers follow the non-Hermitian evolution for longer times before ultimately deviating from it. On the other hand, in the regime of strong monitoring, when the non-Hermitian dynamics display an area law behavior, the role of quantum jumps is less drastic. While we still see strong fluctuations, the average entanglement entropy follows the no-click limit rather closely all the way to the stationary state. From these first results, we can already anticipate that QJs will have a disruptive effect on the volume law phase of the non-Hermitian SSH model. This can be understood by considering that the stochastic noise due to the jumps breaks the PT-symmetry of the non-Hermitian SSH model, which played a major role in protecting the volume law phase in the no-click limit Gal et al. (2023).
@@
to obtain an estimate for the prefactor $c_{\rm eff}$, sometime called effective central charge $c_{\rm eff}$. As we show in the panels (e) and (d) of Fig. (3), this quantity vanishes as a function of both $\gamma$ and $h$, indicating a transition into an area-law phase.
@@
As we see from panel (e), this transition occurs for values of $\gamma\simeq 1$, i.e., still within the volume law phase of the no-click problem (See Fig. 2). Furthermore, the critical point $\gamma_{c}$ decreases weakly for $h<2$, as opposed to the no-click result. On the other hand, we note that the effective central charge is non-monotonous as a function of $h$, vanishing around $h\simeq 2$. This result can be understood from the fact that the band of non-Hermitian Hamiltonian becomes flat around that value Gal et al. (2023). We emphasize that the transition between log and area law scaling could well turn into a crossover upon increasing system sizes through a weak-localization mechanism similar to what was recently discussed in the case of free fermions with $U(1)$ symmetry and projective measurements Poboiko et al. (2023a). Although affected by the inevitable finite-size effects, our results show that the volume law phase of the non-Hermitian SSH is not stable when including quantum jumps. As compared with the phase diagram of the no-click limit, we see that not only does the volume-law turn into a logarithmic phase but also that the emerging area law phase is significantly more extended throughout the phase diagram, which is therefore qualitatively and quantitatively affected by the inclusion of quantum jumps.
@@ VI Statistics of Entanglement Gain and Loss
In the same plots in Fig. 1 (c-d) we also consider the joint probability of observing a change $\Delta S_{\rm qj}$ (or $\delta S_{\rm nH}$) in correspondence of a particular value of entanglement entropy $S$, i.e., $P(\Delta S_{\rm qj},S),\ P(\delta S_{\rm nH},S)$ which are obtained by binning stochastic events (QJs or non-Hermitian evolution) according to the entanglement content of the state they act on. The key idea here is to understand whether QJs impact highly or weakly entangled states more. As we see in Fig. 1 (c-d), the tails of the histogram for the QJs broaden up (at least for $\Delta S_{\rm qj}<0$), indicating that highly entangled states are indeed more fragile and prone to be affected by rare QJs.
To condense the rich information contained in the full statistics of entanglement loss and gain we now focus on the first moment of those distributions, corresponding to the average change to the entanglement entropy due to quantum jumps $\overline{\Delta S}_{\rm qj}(S,\ell)$ and to the non-Hermitian evolution $\overline{\delta S}_{\rm nH}(S,\ell)$. We emphasize that these averages  [^1] With a slight abuse of notation, we indicate the average over the gain/loss statistics with an overbar, as for the average over trajectories. depend on $S$ since they are evaluated over the joint distributions $P(\Delta S_{\rm qj},S)$ and $P(\delta S_{\rm nH},S)$, i.e.
@@
and similarly for $\overline{\delta S}_{\rm nH}(S,\ell)$. In Fig. 7 (a-b), we plot these first moments for the monitored Ising chain (at a particular point of the phase diagram) for different subsystem sizes $\ell$. In panel (a), we see that the change due to non-Hermitian evolution is on average positive, $\overline{\delta S}_{\rm nH}>0$, i.e., it induces an entanglement entropy gain. This gain is substantial for weakly entangled states, then decreases slowly with the entanglement content. In panel (b), we plot the change due to QJs, normalized with respect to the average waiting time $\overline{\tau}$. As shown in Appendix B, this ratio remains well defined upon increasing the system size $L$, despite the waiting time vanishing as $\overline{\tau}\sim 1/L$. This is consistent with the idea that a single quantum jump can affect the entanglement entropy of a quantity $O(1/L)$. From Fig. 7(b), we see that $\overline{\delta S}_{\rm qj}/\overline{\tau}$ is negative and decreases with $S$, i.e. QJs on average induce an entanglement entropy loss. They do so the more the state onto which they act is entangled. Interestingly, both averages develop a subsystem-size $\ell$ dependence above a certain threshold entanglement entropy value. A natural question is how these average entanglement gains and losses are connected to the entanglement dynamics under QJs and its phase transition.
@@ VII.1 Monitored SSH
From this result, we conclude that for the monitored SSH, the logarithmic growth of the entanglement found in Sec. IV arises mainly due to QJs. The picture is that the main effect of increasing the subsystem size is to decrease the slope of $\Delta S_{\rm qj}$, allowing the system to reach a higher value of entanglement entropy. From this point of view, the growth of $S$ should be understood more as a suppression of entanglement loss due to the jumps being less effective in disentangling a large subsystem. In the area law phase (Fig 9 (b)), on the other hand, we see that both contributions are almost independent on $\ell$, and their crossing occurs in a regime of small entanglement entropy, leading to an area-law scaling.
@@ VII.2 Monitored Ising Chain
VII.3 Implications for Free Fermions MIPT
@@
We conclude with a discussion of the implications of our results for the MIPT of free fermions under quantum jumps. The irrelevance of QJs at weak monitoring in the Ising chain is suggestive of a logarithmic phase which could be possibly stable, differently from the SSH case where the jump contributions are likely to saturate at large system sizes through a mechanism similar to the weak-localization correction Poboiko et al. (2023a).
@@ VIII Conclusions
In this work, we studied the role of quantum jumps on the entanglement dynamics of two models of monitored free fermions, the SSH and Ising chains. In both cases, the no-click limit, corresponding to a deterministic non-Hermitian evolution, was studied in detail and revealed a rich pattern of entanglement transitions. The inclusion of QJ has a somewhat different effect in the two models. On one hand, in the SSH case, the inclusion of jumps qualitatively the physics of the no-click limit, washing away the volume-to-area law transition and leaving a sub-volume phase with logarithmic scaling of entanglement entropy, at least for the system sizes considered in this work. On the other hand, for the Ising chain, our results confirm that the effect of stochastic jumps is mainly quantitative, except in specific areas of the phase diagram where deviations from the no-click limit appear.
To better understand the origin of such a different impact of QJs, we have first computed their statistics, particularly their waiting-time distribution. Surprisingly, we have found that at long times, both models acquire a universal Poisson distribution of waiting times, with the same average waiting time $\bar{\tau}$. Then, we introduced an unexplored characterization based on the entanglement gain and loss statistics after or between quantum jumps. The resulting histogram display, particularly for the quantum jump case, has a very broad distribution and a typical value pinned close to zero, suggesting that most QJs are not significantly affecting the entanglement controlled by rare jump events.
@@ Appendix A Statistics of the monitored observable
A different perspective on the role of quantum jumps in the monitored dynamics of our two models can be obtained by looking at the statistics of the monitored observable, in this case, the histogram of the local density. As discussed in Sec. III, this provides information that is not contained in the average dynamics described by the Lindbladian, which only captures the average value (first moment) of the histogram.
To appreciate this point in Fig. 11, we plot for the monitored SSH the statistics of the particle density in the $A$ sublattice for increasing values of the monitoring rate $\gamma$. Given the choice of our initial state, this histogram at time $t=0$ is peaked at $n_{A}=0$, but at long times, we see a redistribution due to the interplay of non-Hermitian dynamics and jumps. For weak monitoring, we see that the long-time limit is an almost flat distribution where all measurement outcomes are almost equally probable, except for the region near $n_{A}=1$, which features a small peak reminiscent of the fact that jumps in the SSH model gets rid of state with empty sites in the $A$ sublattice. For comparison, we show the average density predicted by the non-Hermitian evolution as a dashed line.

### 17 Jan 2024 v2

@@ IV Entanglment Dynamics under Quantum Jumps
In this section we present our results for the QJ dynamics of the SSH and Ising chain model, focusing on the dynamics of the entanglement entropy. Unless specified otherwise, throughout this section we consider an initial product state of the form $|\Psi(0)\rangle=|0,1,0,1,\dots,0,1\rangle$ and open boundary conditions. We fix the hopping along the chains as units of energy and inverse time, $J=1$. We integrate the dynamics with a time-step $dt\propto 1/({4\gamma L})$. Throughout this work we consider a system of size $L=128$.
@@ IV.1 Entanglement Dynamics in the Monitored SSH Model
We start discussing the entanglement dynamics for the monitored SSH. In Fig. 3 (a-c), we show the evolution of the entanglement entropy for a sample of quantum trajectories for increasing values of the monitoring rate $\gamma=0.5,1.5,2.5$, corresponding to the volume-law [panels (a,b)] and area-law [panel (c)] phase of the no-click limit, cf. Fig. 2. For comparison, we also plot the average entanglement entropy (obtained averaging over the different trajectories, according to Eq. (12)) as well as the value obtained in the non-Hermitian no-click limit. We see that, for weak monitoring, when the no-click dynamics displays a linear growth corresponding to the volume law phase, the stochastic dynamics due to quantum jumps suppress the entanglement very effectively and bring the system to a stationary state which is less entangled than in the no-click limit. In this regime, an agreement with pure non-Hermitian evolution is present only for short time scales. We also notice that for weak monitoring, trajectory-to-trajectory fluctuations are rather broad, and certain outliers follow the non-Hermitian evolution for longer times before ultimately deviating from it. On the other hand, in the regime of strong monitoring, when the non-Hermitian dynamics display an area law behavior, the role of quantum jumps is less drastic. While we still see strong fluctuations, the average entanglement entropy follows the no-click limit rather closely all the way to the stationary state. From these first results, we can already anticipate that QJs will have a disruptive effect on the volume law phase of the non-Hermitian SSH model. This can be understood by considering that the stochastic noise due to the jumps breaks the PT-symmetry of the non-Hermitian SSH model, which played a major role in protecting the volume law phase in the no-click limit Gal et al. (2023).
@@
to obtain an estimate for the prefactor $c_{\rm eff}$, sometime called effective central charge $c_{\rm eff}$. As we show in the panels (e) and (f) of Fig. (3), this quantity vanishes as a function of both $\gamma$ and $h$, indicating a transition into an area-law phase.
@@
As we see from panel (e), this transition occurs for values of $\gamma\simeq 1$, i.e., still within the volume law phase of the no-click problem (See Fig. 2). Furthermore, the critical point $\gamma_{c}$ decreases weakly for $h<2$, as opposed to the no-click result. On the other hand, we note that the effective central charge is non-monotonous as a function of $h$, vanishing around $h\simeq 2$. This result can be understood from the fact that the quasiparticle band of the non-Hermitian Hamiltonian becomes flat around that value Gal et al. (2023) and, as a consequence, the QJ dynamics at short-intermediate times slows down. We emphasize that the entanglement transition between log and area law scaling discussed above could well turn into a crossover upon increasing system sizes through a weak-localization mechanism similar to what was recently discussed in the case of free fermions with $U(1)$ symmetry and projective measurements Poboiko et al. (2023a). Although affected by the inevitable finite-size effects, our results show that the volume law phase of the non-Hermitian SSH is not stable when including quantum jumps. As compared with the phase diagram of the no-click limit, we see that not only does the volume-law turn into a logarithmic phase but also that the emerging area law phase is significantly more extended throughout the phase diagram, which is therefore qualitatively and quantitatively affected by the inclusion of quantum jumps.
@@ VI Statistics of Entanglement Gain and Loss
In the same plots in Fig. 1 (c-d) we also consider the joint probability of observing a change $\Delta S_{\rm qj}$ (or $\delta S_{\rm nH}$) in correspondence of a particular value of entanglement entropy $S$, i.e., $P(\Delta S_{\rm qj},S),\ P(\delta S_{\rm nH},S)$ which are obtained by binning stochastic events (QJs or non-Hermitian evolution) according to the entanglement content of the state they act on. The key idea here is to understand whether QJs impact differently states which are highly or weakly entangled. As we see in Fig. 1 (c-d), the tails of the histogram for the QJs broaden up (at least for $\Delta S_{\rm qj}<0$), indicating that highly entangled states are indeed more fragile and prone to be affected by rare QJs.
To condense the rich information contained in the full statistics of entanglement changes we now focus on the first moment of those distributions, corresponding to the average change to the entanglement entropy due to quantum jumps $\overline{\Delta S}_{\rm qj}(S,\ell)$ and to the non-Hermitian evolution $\overline{\delta S}_{\rm nH}(S,\ell)$. We emphasize that these averages  [^1] With a slight abuse of notation, we indicate the average over the gain/loss statistics with an overbar, as for the average over trajectories. depend on the average entanglement entropy $S$ since they are evaluated over the joint distributions $P(\Delta S_{\rm qj},S)$ and $P(\delta S_{\rm nH},S)$, i.e.
@@
and similarly for $\overline{\delta S}_{\rm nH}(S,\ell)$. In Fig. 7 (a-b), we plot these first moments for the monitored Ising chain (at a particular point of the phase diagram) for different subsystem sizes $\ell$. In panel (a), we see that the change due to non-Hermitian evolution is on average positive, $\overline{\delta S}_{\rm nH}>0$, i.e., it induces an entanglement entropy gain. This gain is substantial for weakly entangled states, then decreases slowly with the entanglement content. In panel (b), we plot the change due to QJs, normalized with respect to the average waiting time $\overline{\tau}$. As shown in Appendix B, this ratio remains well defined upon increasing the system size $L$, despite the waiting time vanishing as $\overline{\tau}\sim 1/L$. This is consistent with the idea that a single quantum jump can affect the entanglement entropy of a quantity $O(1/L)$. From Fig. 7(b), we see that $\overline{\delta S}_{\rm qj}/\overline{\tau}$ is negative and decreases with $S$, i.e. QJs on average induce an entanglement entropy loss. They do so the more the state onto which they act is entangled. Interestingly, both averages develop a subsystem-size $\ell$ dependence above a certain threshold entanglement entropy value. A natural question that we adress next is how these average entanglement gains and losses are connected to the entanglement dynamics under QJs and its phase transition.
@@ VII.1 Monitored SSH
From this result, we conclude that for the monitored SSH, the logarithmic growth of the entanglement found in Sec. IV arises mainly due to QJs. The picture is that the main effect of increasing the subsystem size is to decrease the slope of $\Delta S_{\rm qj}$, i.e. decrease the rate at which QJs reduce entanglement thus allowing the system to reach a higher value of entanglement entropy. From this point of view, the growth of $S$ should be understood more as a suppression of entanglement loss, due to the jumps being less effective in disentangling a large subsystem. For large monitoring (Fig 9 (b)), on the other hand, we see that both contributions are almost independent on $\ell$, and their crossing occurs in a regime of small entanglement entropy, leading to an area-law scaling.
@@ VII.2 Monitored Ising Chain
VII.3 Discussion and Implications for Free Fermions MIPT
@@
We conclude with a discussion of the implications of our results for the MIPT of free fermions under quantum jumps. The irrelevance of QJs at weak monitoring in the Ising chain points towards an important role of the non-Hermitian Hamiltonian in this regime. This suggests that the logarithmic phase of entanglement entropy under QJs could be possibly stable in the large system size, as in the no-click case Turkeshi and Schiró (2023). This is consistent with the replica field theory prediction for noisy Majorana fermions under QSD Fava et al. (2023), although we emphasize the differences in measurement protocol and unitary dynamics between the two cases. In particular based on our numerical results we cannot establish whether also in the Ising chain the entanglement scaling is $\ln^{2}(\ell)$. On the other hand our results for the SSH case points toward a strong effect of QJs at weak monitoring. As we have shown the logarithmic phase found numerically is mainly due to QJs. Their contribution is likely to saturate at large system sizes through a mechanism similar to the weak-localization correction Poboiko et al. (2023a).
@@ VIII Conclusions
In this work, we studied the role of quantum jumps on the entanglement dynamics of two models of monitored free fermions, the SSH and Ising chains. In both cases, the no-click limit, corresponding to a deterministic non-Hermitian evolution, was studied in detail and revealed a rich pattern of entanglement transitions. The inclusion of QJ has a somewhat different effect in the two models. On one hand, in the SSH case, the inclusion of jumps qualitatively changes the physics of the no-click limit, washing away the volume-to-area law transition and leaving a sub-volume phase with logarithmic scaling of entanglement entropy, at least for the system sizes considered in this work. On the other hand, for the Ising chain, our results confirm that the effect of stochastic jumps is mainly quantitative, except in specific areas of the phase diagram where deviations from the no-click limit appear.
To better understand the origin of such a different impact of QJs, we have first computed their statistics, particularly their waiting-time distribution. Surprisingly, we have found that at long times, both models acquire a universal Poisson distribution of waiting times, with the same average waiting time $\bar{\tau}$. Then, we introduced an unexplored characterization based on the entanglement gain and loss statistics after or between quantum jumps. The resulting histogram display, particularly for the quantum jump case, a very broad distribution and a typical value pinned close to zero, suggesting that most QJs are not significantly affecting the entanglement controlled by rare jump events.
@@ Appendix A Statistics of the monitored observable
A different perspective on the role of quantum jumps in the monitored dynamics of our two models can be obtained by looking at the statistics of the monitored observable, in this case, the histogram of the local density. As discussed in Sec. III of the main text, this provides information that is not contained in the average dynamics described by the Lindbladian, which only captures the average value (first moment) of the histogram.
To appreciate this point in Fig. 11 we plot for the monitored SSH the statistics of the particle density in the $A$ sublattice for increasing values of the monitoring rate $\gamma$. Given the choice of our initial state, this histogram at time $t=0$ is peaked at $n_{A}=0$, but at long times we see a redistribution due to the interplay of non-Hermitian dynamics and jumps. For weak monitoring, we see that the long-time limit is an almost flat distribution where all measurement outcomes are almost equally probable, except for the region near $n_{A}=1$, which features a small peak reminiscent of the fact that jumps in the SSH model gets rid of state with empty sites in the $A$ sublattice. For comparison, we show the average density predicted by the non-Hermitian evolution as a dashed line.